# Deep Evidential Learning for Dose Prediction

### v1

@@ 4 Results
Table 1: Table collecting various measures of uncertainty-error correlations. Abbreviations: MAE = mean absolute error, $U_{avg}$ = mean uncertainty value, $r_{s}(\overline{U},\overline{D_{e}})$ = Spearman’s coefficient between patient-averaged uncertainty and error distributions, $r_{s}(U,D_{e})$ = Spearman’s coefficient between uncertainty and error distributions, $M.I.(U,D_{e})$ = mutual information between uncertainty and error distributions, $D_{e}$ = prediction error, $(U_{e},U_{a})$ = (epistemic uncertainty, aleatoric uncertainy).
@@ 4.1 Variation of prediction error with uncertainty threshold
The results in Table 1 indicated that, at least compared to Monte-Carlo Dropout and Deep Ensemble methods, Deep Evidential Learning model appeared to yield spectra of uncertainty values which were strongly correlated with prediction error, indicating that it could enact a feasible uncertainty estimation framework that helps to discover and highlight potential regions of ignorance within the deep learning model.
To further scrutinize the difference among the various frameworks, we examined how error distributions varied with changing threshold values of uncertainty. Since the error distributions were generically found to be quite skewed, we used the median of the error distribution as its characterizing parameter. Fig. 3 and 4 below reveal how median error changed with increasing levels of uncertainty thresholds for the different models. Assuming that uncertainty values correlate positively with prediction error implies that one expects a monotonically increasing curve in such a plot, a trend that was indeed manifest for these models as shown in Fig. 3 and 4.
@@ 4.2 Probing noise-sensitivity of uncertainty measures
A prominent feature of Deep Evidential Learning is a clear principled approach towards distinguishing between aleatoric and epistemic uncertainties. They can be defined precisely using the higher-order PDF $\mathcal{P}$ as expressed in eqn. (5). Traditionally, the definition of aleatoric uncertainty is commonly taken to imply that it can be understood to capture any inherent random noise in the data, in contrast to epistemic uncertainty which presumably captures uncertainty originating from model-suitability and data-sufficiency (see e.g. [32, 36]).
@@ 4.3.1 Uncertainty heatmaps
ROI $\overline{U}_{alea}$ (Gy2) $\overline{U}_{epis}$ (Gy2) $r_{s}(U_{alea},D_{e})$ $r_{s}(U_{epis},D_{e})$ $\overline{D}_{e}=\overline{|D_{p}-D_{GT}|}$ (Gy)
@@
Table 2: Table collecting various ROIs, their associated uncertainty ($U_{alea},U_{epis}$) values together with the errors in predicted dose $D_{e}$. These values are averaged over the entire testing dataset. $r_{s}(U_{alea},D_{e})$ and $r_{s}(U_{epis},D_{e})$ denote the Spearman’s coefficients between each uncertainty type and the corresponding dose prediction error $D_{e}$.
At the level of each ROI, apart from the larynx, $r_{s}(U_{epis},D_{e})\gtrapprox 0.7$ (1 s.f.) for all other OAR and PTV regions. Coupled with the error-uncertainty threshold study in Sec. 4.1 and the noise-sensitivity test in Sec. 4.2, this tapestry of results equipped us with the basis for interpreting elevated heatmap regions of epistemic uncertainty as indicators of potential regions of prediction errors and those of aleatoric uncertainty as correlated with inherent data noise.
In Fig. 6, we display an illustrative set of axial CT images portrayed alongside their corresponding uncertainty heatmaps. In principle, the aleatoric uncertainty probes the level of inherent data noise, while the epistemic uncertainty is more strongly correlated with the elevated regions in the (leftmost) dose prediction error heatmap. Thus, we expect the latter to be more heterogeneous which was indeed evident in our survey of uncertainty heatmaps across the testing dataset.
Figure 6: Axial CT slices with their model prediction error and their corresponding uncertainty heatmaps. In each row, the leftmost image shows the mean absolute dose prediction error (MAE) heatmap, the center and rightmost images showing the aleatoric and epistemic uncertainty distributions respectively.
@@ 4.3.2 DVH with confidence bands
where conditional expectations such as $\mathbb{E}\left[Y|X\right]$ are random variables themselves. To apply (12) appropriately to our context, we can identify $Y\sim\mu_{p}$, and $X$ to collectively represent all model-dependent and data-dependent variables. The relation (11) then follows directly from (12) by virtue of the definitions of $\sigma^{2},\mu$. In the following, we furnish a contextual proof of this result taking into account our framework explicitly.iiiiiiiiiOur style of proof is similar in spirit to that presented in [41] for a class of Deep Ensemble models that admit interpretations as Gaussian mixture models, and of which outputs are engineered to be $\mu,\sigma$. We begin with the definition
@@
In Fig. 7, we plot examples of individual patient’s DVH for various PTV and OAR regions, each equipped with a $95\%$ confidence interval. A visible feature of these bands is that they tend to contain the more extreme deviations of each groundtruth DVH from its corresponding predicted one along their peripheral edges. This is consistent with the order-of-magnitude estimate of $MAE/\delta D_{p}\sim 1.7$ as drawn from numerical values of the mean error and uncertainties in Table 1. Thickness and various features of these DVH confidence bands for the OARs and PTV regions can be used to characterize the extent of reliability of various dose prediction models, differentiating among them in this aspect.
Figure 7: A set of examples of DVH for spinal cord, esophagus, right parotid and larynx each equipped with confidence bands. Dashed curves are the bounds for the $95\%$ intervals defined by ±$\pm 1.96\,\delta D_{p}\equiv\pm 1.96\sqrt{U^{2}_{alea}+U^{2}_{epis}}$. Red and blue curves pertain to the ground truth and model prediction respectively. In each diagram, $\overline{\delta D}$ refers to the standard error in dose in units of Gy obtained by averaging $\delta D_{p}$ over all voxels of the ROI in the patient.
@@ 5 Discussion
We have refined and applied Deep Evidential Learning framework to radiotherapy dose prediction, our most prominent finding being that the uncertainty estimates obtained from model training inherited strong correlations with prediction errors. In the following, we conclude with a summary list of key findings accompanied by a discussion of various limitations of our work.
5.1 Summary list of key findings
  1. (a)
    The original loss function proposed in [29] did not work in this particular context of dose prediction. We found that regularizing the final layer with a sigmoid function and adding a mean-squared-error term led to a stable and effective implementation of Deep Evidential Learning.
  2. (b)
    Epistemic uncertainty ($U_{epis}$) in Deep Evidential Learning was highly correlated with prediction errors. Correlation indices (mutual information, Spearman’s coefficients) were comparable or higher than those for MC Dropout and Deep Ensemble methods (see Table 1).
  3. (c)
    The median error varied with uncertainty threshold much more linearly for $U_{epis}$ as compared to MC Dropout and Deep Ensemble, indicative of a more uniformly calibrated sensitivity to model errors (see Fig. 3, 4).
  4. (d)
    Aleatoric uncertainty $(U_{alea})$ demonstrated a more significant shift in its empirical CDF upon addition of Gaussian noise to CT intensity distribution compared to $U_{epis}$ (see Fig. 5). It also displayed visibly weaker correlation with prediction errors relative to $U_{epis}$. These traits of $U_{alea},U_{epis}$ appeared to be supportive of their conventional interpretations as reflecting data noise and model-related uncertainties respectively.
  5. (e)
    We demonstrated how Eve’s law of total variance enables one to express the predictive variance in terms of $U_{alea},U_{epis}$, and used this result to construct confidence intervals for DVH (see Fig. 8) as an application that can be further used to characterize reliability of the model.
@@ 5.2 Limitations and Future Directions
In the following, we outline several limitations of our work together with corresponding suggestions for future directions.
  1. (i)
    Although the simple 3D U-Net backbone architecture led to efficient and fast convergence, it would be interesting to study how more complicated network structures like the transformers-based models of [17, 18] perform after modifying their final layers so that there are four output heads corresponding to the parameters of $\mathcal{P}$ in Deep Evidential Learning. This will provide a more extensive study of the degree to which a low MAE can be attained while not compromising the uncertainty-error correlation.
  2. (ii)
@@
  3. (iii)
    We had used the same regularization term $\mathcal{L}_{reg}=|y-\gamma|(2\nu+\alpha)$ in the loss function following [29], which is independent of the $\beta$ parameter of the prior distribution $\mathcal{P}$ in eqn. (4). It would be interesting to explore if there are other alternatives which express full dependence on all four parameters of $\mathcal{P}$. In an analogous framework for classification, Sensoy et al. in [30] showed that the regularization term measuring the Kullback-Leibler divergence from an uninformative prior (e.g. uniform distribution) worked well for classification problems, and notably, the authors of [29] had attempted to find the corresponding version for regression yet without success.

### v2

@@ 4 Results
Table 1: Table collecting various measures of uncertainty-error correlations. Abbreviations: MAE = mean absolute error, $U_{avg}$ = mean uncertainty value, $r_{s}(\overline{U},\overline{D_{e}})$ = Spearman’s coefficient between patient-averaged uncertainty and error distributions, $r_{s}(U,D_{e})$ = Spearman’s coefficient between uncertainty and error distributions, $M.I.(U,D_{e})$ = mutual information between uncertainty and error distributions, $D_{e}$ = prediction error, $(U_{e},U_{a})$ = (epistemic uncertainty, aleatoric uncertainty).
@@ 4.1 Variation of prediction error with uncertainty threshold
The results in Table 1 indicated that Deep Evidential Learning model yielded spectra of uncertainty values which were more correlated with prediction error, relative to MC Dropout and Deep Ensemble methods. To further scrutinize the difference among the various frameworks, we examined how error distributions varied with changing threshold values of uncertainty. Since the error distributions were generically found to be quite skewed, we used the median of the error distribution as its characterizing parameter. Fig. 3 and 4 below reveal how median error changed with increasing levels of uncertainty thresholds for the different models. Assuming that uncertainty values correlate positively with prediction error implies that one expects a monotonically increasing curve in such a plot, a trend that was indeed manifest for these models as shown in Fig. 3 and 4.
@@ 4.2 Probing noise-sensitivity of uncertainty measures
A prominent feature of Deep Evidential Learning is a clear principled approach towards distinguishing between aleatoric and epistemic uncertainties. They can be defined precisely using the higher-order PDF $\mathcal{P}$ as expressed in eqn. (5). Traditionally, aleatoric uncertainty is commonly interpreted as arising from inherent random noise in the data, in contrast to epistemic uncertainty which presumably captures uncertainty originating from model suitability and data sufficiency (see e.g. [32, 36]).
@@ 4.3.1 Uncertainty heatmaps
ROI $\overline{U}_{alea}$ (Gy2) $\overline{U}_{epis}$ (Gy2) $r_{s}(U_{alea},D_{e})$ $r_{s}(U_{epis},D_{e})$ $\overline{D}_{e}$ (Gy)
@@
Table 2: Table collecting various ROIs, their associated uncertainty ($U_{alea},U_{epis}$) values together with the errors in predicted dose $D_{e}=|D_{p}-D_{GT}|$. These values were averaged over the entire testing dataset. $r_{s}(U_{alea},D_{e})$ and $r_{s}(U_{epis},D_{e})$ denote the Spearman’s coefficients between each uncertainty type and the corresponding dose prediction error $D_{e}$.
At the level of each ROI, apart from the larynx, $r_{s}(U_{epis},D_{e})\gtrapprox 0.7$ (1 s.f.) for all other OAR and PTV regions. Coupled with the error-uncertainty threshold study in Sec. 4.1 and the noise-sensitivity test in Sec. 4.2, this tapestry of results equips us with the basis for interpreting elevated heatmap regions of epistemic uncertainty as indicators of potential regions of prediction errors.
In Fig. 6, we display an illustrative set of axial CT images portrayed alongside their corresponding uncertainty heatmaps. In principle, the aleatoric uncertainty probes the level of inherent data noise, while the epistemic uncertainty is more strongly correlated with the elevated regions in the (leftmost) dose prediction error heatmap. Indeed, in Fig. 6, we observe that the relative intensity map of epistemic uncertainty tends to parallel the corresponding error heatmap more sensitively.
Figure 6: Axial CT slices depicting model prediction errors and their uncertainty heatmaps. Each row corresponds to an example patient in the testing dataset of [14]; panels (a) and (d) show the mean absolute dose prediction error heatmap; panels (b) and (e) show the aleatoric uncertainty distribution, while panels (c) and (f) depict the epistemic uncertainty distribution (normalized within each image).
@@ 4.3.2 DVH with confidence bands
where conditional expectations such as $\mathbb{E}\left[Y|X\right]$ are random variables themselves. To apply (12) appropriately to our context, we can identify $Y\sim\mu_{p}$, and $X$ to collectively represent all model-dependent and data-dependent variables. The relation (11) then follows directly from (12) by virtue of the definitions of $\sigma^{2},\mu$. In the following, we furnish a contextual proof of this result taking into account our framework explicitly.vvvOur style of proof is similar in spirit to that presented in [49] for a class of Deep Ensemble models that admit interpretations as Gaussian mixture models, and of which outputs are engineered to be $\mu,\sigma$. We begin with the definition
@@
In Fig. 7, we plot examples of individual patient’s DVH for various PTV and OAR regions, each equipped with a $95\%$ confidence interval. In each diagram of Fig. 7, $\overline{\delta D}$ refers to the standard error in dose in units of Gy obtained by averaging $\delta D_{p}$ over all voxels of the ROI in the patient whose ID shown is defined in the dataset of [14]. A visible feature of these bands is that they tend to contain the more extreme deviations of each groundtruth DVH from its corresponding predicted one along their peripheral edges. This is consistent with the order-of-magnitude estimate of $MAE/\delta D_{p}\sim 1.7$ as drawn from numerical values of the mean error and uncertainties in Table 1. Thickness and various features of these DVH confidence bands for the OARs and PTV regions can be used to characterize the extent of reliability of various dose prediction models, differentiating among them in this aspect.
Figure 7: A set of examples of DVH for spinal cord, esophagus, right parotid and larynx each equipped with confidence bands. Dashed curves are the bounds for the $95\%$ intervals defined by ±$\pm 1.96\,\delta D_{p}\equiv\pm 1.96\sqrt{U^{2}_{alea}+U^{2}_{epis}}$. Red and blue curves pertain to the ground truth and model prediction respectively.
Fig. 8 presents an illustrative DVH containing a number of organs-at-risk and target regions for an individual patient in the dataset of [14]. Such a confidence band-enhanced DVH can be employed in the radiotherapy clinic to enable a more reliable interpretation of the model-predicted DVH.
@@ 5 Discussion
We have refined and applied Deep Evidential Learning framework to radiotherapy dose prediction, our most prominent finding being that the uncertainty estimates obtained from model training inherited strong correlations with prediction errors. In the following, we conclude with a summary of key findings accompanied by a discussion of various limitations of our work.
5.1 Summary of key findings
The original loss function proposed in [29] did not work in this particular context of dose prediction. We found that regularizing the final layer with a sigmoid function and adding a mean-squared-error term led to a stable and effective implementation of Deep Evidential Learning. Upon completion of model training, we found that the model inherited uncertainty estimates at a more granular level compared to conventional MC Dropout and Deep Ensemble methods. Epistemic uncertainty ($U_{epis}$) in Deep Evidential Learning was highly correlated with prediction errors. Correlation indices (mutual information, Spearman’s coefficients) were comparable or higher than those for MC Dropout and Deep Ensemble methods (see Table 1). Aleatoric uncertainty $(U_{alea})$ demonstrated a more significant shift in its empirical CDF upon addition of Gaussian noise to CT intensity distribution compared to $U_{epis}$ (see Fig. 5). It also displayed visibly weaker correlation with prediction errors relative to $U_{epis}$. These traits of $U_{alea},U_{epis}$ appeared to be supportive of their conventional interpretations as reflecting data noise and model-related uncertainties respectively.
Another aspect of error-uncertainty relationship that appeared to distinguish Deep Evidential Learning was that the median error varied with uncertainty threshold much more linearly for $U_{epis}$ as compared to MC Dropout and Deep Ensemble methods, indicative of a more uniformly calibrated sensitivity to model errors (see Fig. 3, 4). This is compatible with the fact that the uncertainty heatmaps generated by the model were found to be highly effective in identifying potential regions of model’s inaccuracies.
5.2 Towards incorporating our models in the radiotherapy clinic
An essential plan evaluation tool for the radiation oncologist is the DVH that characterizes a dose distribution. We demonstrated how Eve’s law of total variance enables one to express the predictive variance in terms of $U_{alea},U_{epis}$, and used this result to construct confidence intervals for DVH (see Fig. 8). This technique can be harnessed by the radiotherapy treatment planning team to enable a more statistically informed interpretation of DVH in deep learning-guided treatment planning. In the Knowledge-Based-Planning pipeline, dose prediction models can be used as inputs to some dose mimicking model [7] that produces deliverable treatments via optimization of a set of objective functions [25, 26]. The probabilistic dose distributions generated by our models can serve as robust inputs to the objective functions of [25, 26] to generate details of deliverable radiotherapy treatments. In this regard, it would be interesting to explore how the uncertainty estimates generated by our models translate to those of related quantities in the final treatment plan, such as the positions and motion dynamics of multileaf collimators and other machine parameters. This would pave the way for completing an uncertainty-aware AI-guided pipeline for radiotherapy treatment planning.
@@ 5.3 Limitations and Future Directions
Finally, we outline several limitations of our work together with corresponding suggestions for future directions. Although the simple 3D U-Net backbone architecture led to efficient and fast convergence, it would be interesting to study how more complicated network structures like the transformers-based models of [17, 18] perform after modifying their final layers so that there are four output heads corresponding to the parameters of $\mathcal{P}$ in Deep Evidential Learning.
As a simple explicit example, a U-Net-like transformer model (Swin-UNet) was proposed in [51] and [52] for segmentation tasks. To enable Deep Evidential Learning on such backbone models, one can append a final 4-channel pointwise convolution layer to these transformer block-based architectures, which will yield the four parameters of the prior normal-inverse-gamma distribution in eqn. (4) as model outputs. Model training should employ the maximum likelihood loss function of [29] or the refined version that we have proposed here in eqn. (10). More generally, this simple minimal prescription of appending a final layer of 4-channel pointwise convolution layer applies for other transformer-based vision models adapted for segmentation. If higher model complexity is needed, this final layer can be preceded by more multi-channel pointwise convolution layers, with other essential traits of transformer models left intact, e.g. multi-head self-attention modules, shifted window-based attention mechanism, etc. Details of these inner layers are unaffected and do not require additional modifications when encapsulating the transformer-based vision model within the framework of Deep Evidential Learning.
Enabling Deep Evidential Learning on other backbone model structures will provide a more extensive study of the degree to which a low MAE can be attained while not compromising the uncertainty-error correlation. Indeed it will be interesting to explore how our refined Deep Evidential Learning model works for other regression tasks beyond the context of dose prediction, such as traffic forecasting in telecommunication networks in [53, 54]. It would be interesting to pursue whether the variety of model architectures examined in [53, 54] can be extended to uncertainty-aware ones using our techniques.
@@
We had used the same regularization term $\mathcal{L}_{reg}=\left|\log\left(\frac{y}{1-y}\right)-\gamma\right|(2\nu+\alpha)$ in the loss function following [29], which is independent of the $\beta$ parameter of the prior distribution $\mathcal{P}$ in eqn. (4). It would be interesting to explore if there are other alternatives which express full dependence on all four parameters of $\mathcal{P}$. In an analogous framework for classification, Sensoy et al. in [30] showed that the regularization term measuring the Kullback-Leibler divergence from an uninformative prior (e.g. uniform distribution) worked well for classification problems, and notably, the authors of [29] had attempted to find the corresponding version for regression yet without success.
Fundamentally, our modeling of dose prediction was based on the simplifying assumption that the patient’s CT images and the oncologist’s specification of the target regions are sufficient inputs for predicting the pareto-optimal dose distribution. In reality, effectiveness of radiation is often sensitive to a more complex web of biological factors. Nonetheless, we hope that our work can be a good starting point towards precision radiation oncology. Towards this goal, a pertinent future direction would be to expand the set of model’s input features to include genetic, molecular and other unique clinical characteristics of each patient, beyond just using medical images.
@@ 6 Conclusion
Declaration of competing interest
All authors have no conflict of interest to declare.
Acknowledgments
H.S.Tan wishes to acknowledge that the primary development of this work was conducted while he was affiliated with University of Pennsylvania, Perelman School of Medicine, Department of Radiation Oncology, to whom he is grateful for support.